# Demystifying optimized prompts in language models

### v1

@@ Abstract
Modern language models (LMs) are not robust to out-of-distribution inputs. Machine generated (“optimized”) prompts can be used to modulate LM outputs and induce specific behaviors while appearing completely uninterpretable. In this work, we investigate the composition of optimized prompts, as well as the mechanisms by which LMs parse and build predictions from optimized prompts. We find that optimized prompts primarily consist of punctuation and noun tokens which are more rare in the training data. Internally, optimized prompts are clearly distinguishable from natural language counterparts based on sparse subsets of the model’s activations. Across various families of instruction-tuned models, optimized prompts follow a similar path in how their representations form through the network.   [^1] Our code and models will be openly available upon publication.
@@ 1 Introduction
Language models (LMs) (Grattafiori et al., 2024; Biderman et al., 2023a; Team et al., 2024; Abdin et al., 2024) are trained on large amounts of filtered internet data (Gao et al., 2020; Raffel et al., 2020; Penedo et al., 2024; Soldaini et al., 2024), which consist primarily of interpretable natural language text. Recent work has found that these models are sensitive to machine-generated optimized prompts, which, although seemingly uninterpretable, can be used to elicit targeted behaviors (Shin et al., 2020; Wen et al., 2023; Zou et al., 2023b; Melamed et al., 2024).
@@ 1.1 Our contributions
When comparing tokens in both optimized and natural language prompts to the pre-training corpus, we find that the majority of tokens in optimized prompts are more rare than their natural language counterparts. Furthermore, the token distribution of optimized prompts visibly deviates from standard Zipfian behavior; see Section 5.
@@ 3 Experimental setup
We focus our work on transformer decoder (Vaswani et al., 2017; Radford and Narasimhan, 2018) models. We use the Tiny Stories Eldan and Li (2023) dataset, which consists of synthetically generated stories meant to be understandable by a three year old child. We train a transformer decoder language model based on the GPT-NeoX Black et al. (2022); Biderman et al. (2023a) architecture; see Appendix A for full training details. Originally, the model uses a Byte-pair encoding (BPE) tokenizer (Sennrich et al., 2016), which results in optimized prompts that include several nonsensical characters and subwords Melamed et al. (2024); Cherepanova and Zou (2024). Because we wish to better understand which specific words appear in optimized prompts, we train a new word-level tokenizer over the Tiny Stories corpus. Using word-level tokenization allows us to better interpret optimized prompts, since we do not need to extrapolate meaning from sub-word tokens and can directly evaluate each word in the prompt individually.
@@
In addition to the word-level Tiny Stories model, we optimize prompts using 17 open models from various model families, including both base models, and their instruction-tuned variants which have been aligned for chat purposes. We use a variety of datasets for the optimization including Alpaca (Taori et al., 2023), WikiText-103 (Salesforce, 2021), OpenHermes-2.5 (Teknium, 2023), and Dolly-15k (Conover et al., 2023); see Appendix A for further details.
@@
$d_{KL}({\boldsymbol{p}}^{*}||{\boldsymbol{p}})=\frac{1}{n}\sum_{i=1}^{n}\log(\mathbb{P}_{\mathrm{LM}}({\boldsymbol{d}}_{i}|{\boldsymbol{p}}^{*}))\\ -\log(\mathbb{P}_{\mathrm{LM}}({\boldsymbol{d}}_{i}|{\boldsymbol{p}}))$ (1)
@@ 4 Optimized prompts consist of specific influential tokens
$s_{i}=d_{KL}({\boldsymbol{p}}||{\boldsymbol{p}}_{-i})$ (2)
where ${\boldsymbol{p}}_{-i}=[p_{1},...,p_{i-1},p_{i+1},...,p_{k}]$ is the prompt with token $i$ removed and $d_{KL}$ is defined in Equation 1. A larger influence score indicates that removing token $i$ causes a greater deviation from the functional behavior of the original prompt.
@@ 5.2 Optimized prompts rely on tokens that are rare in the training data
Figure 5: Sparse linear classifier probe accuracy on top-$k$ features of the model output at each layer. We identify top features using Equation 4, and train a linear classifier to discriminate between optimized and natural language prompts. The first column compares optimized and natural language prompts. The second and third columns show a baseline comparison of natural language prompts vs other natural language prompts, and of natural language prompts vs random prompts, respectively. We refer the reader to Appendix D for the full results on all models.
@@ 6.1 Sparse probing for optimized prompts
Figure 6: Average top-10 token prediction overlap. The overlap is computed via Equation 5. Overall, the importance of the top identified features appears to be model dependent, and it is not necessarily the case that the optimized prompts rely more heavily on these features. However, it is clear that there are certain layers which influence the output more (specifically the first and last layers); see Appendix D for the full results.
@@
We then train a logistic regression classifier using the top identified features with varying levels of sparsity. Overall, we find that the classifier is able to discriminate ground truth and optimized examples with high accuracy, even with high levels of sparsity; see Figure 5. This is in contrast to the near-random accuracy found in the baseline evaluations which compare original natural language prompts to randomly generated prompts and other natural language prompts. It is important to note that even though for optimized prompts, models clearly contain unique representations which are easily distinguishable from natural prompts, these representations still generate functionally similar outputs.
@@ 6.2 Do optimized prompts rely on a distinct subspace?
Given that we can classify prompts based on distinct features in the activations, we test the importance of these features in eliciting desired output from optimized prompts. Prior work finds that optimized prompt are more sensitive to discrete perturbations (Ishibashi et al., 2023; Melamed et al., 2024; Cherepanova and Zou, 2024), and we hypothesize that this sensitivity is also present in the model’s internal representations. In order to verify our hypothesis, we perform causal intervention and zero-out the top features identified layer-wise via Equation 4, and then measure the top-10 token overlap for each (original, optimized) prompt pair.
@@
We find that intervening on top features for both natural and optimized prompts has a pronounced effect when compared to the baseline of intervening on random features. However, surprisingly, our experiment contradicts the hypothesis. As shown in Figure 6, for the majority of layers in each model, there is not a large difference between the effect of ablations of top-$k$ features on natural language prompts vs optimized prompts. This means that although optimized prompts may be more sensitive to discrete token-level perturbations than natural language prompts, they do not necessarily share this sensitivity when evaluated from the perspective of the model’s internal representations. We do note that for some models, there are specific layers (first and final) which induce a more pronounced change on the output when top features are ablated.
@@ 6.3 How do LMs build predictions from optimized prompts?
Figure 7: Layer-wise KL divergence. The layer-wise KL divergence is computed as described in Section 6.3. We find that instruction-tuned models follow a similar path; see Appendix D for the full results on all models.
@@ 7 Discussion
Our work analyzes the mechanisms and ways in which language models parse and interpret discretely optimized prompts. We find that optimized prompts consist primarily of punctuation and noun tokens which are, on average, more rare in the training data. Through sparse probing, we are able to classify optimized prompts and their natural language counterparts with high accuracy. Furthermore, when ablating neurons from model layers, the effectiveness of optimized prompts does not drop in a significant way compared to their natural language counterparts.
One possible application of our analysis is to identify optimized “jailbreak” prompts before these prompts are even fully processed by the model. For example, one can train a simple linear classifier on a set of optimized prompts and natural language, and efficiently alert the model provider if a user is inputting suspicious optimized prompts based on the classifier at intermediate layer.
@@
Due to computational constraints, we limit our analysis to models up to 3 billion parameters, as the discrete optimization process requires significant GPU memory and compute. Nevertheless, we test 17 different models. Future work can consider analyzing larger models and investigating trends in optimized prompts across a wider array of model sizes. This will also enable research on scaling effects for prompt optimization.
Acknowledgments
We thank Enric Boix-Adserà for many helpful discussions. This research was developed in part with funding from NSF under grant 2127207.
@@ A.2 Optimization and data setup
For all models, we perform the evil twins optimization procedure for 500 steps, with early stopping if $d_{KL}({\boldsymbol{p}}^{*}||{\boldsymbol{p}})\leq 5.0$. In total, we obtain 5000 unique prompts for the word-stories model, and 1200 unique prompts for the open models (from the four aforementioned prompt datasets), with 300 prompts from each dataset. We then filter all final optimized prompts such that $d_{KL}({\boldsymbol{p}}^{*}||{\boldsymbol{p}})\leq 10.0$. Table 2 displays the number of prompts that were optimized for each model after filtering.
@@ Appendix D Experimental results on all models
We report the full results on the remainder of the 17 tested models for the probing classifier experiment, the feature ablation experiment, and the layer-wise KL divergence experiment. Figure 10 displays the results for all model suites on the sparse probing experiment. Figure 11 displays the results for all model suites on the intervention experiment. Figure 12 displays the results for all model suites on the layer-wise KL divergence experiment.

### v2

@@ Abstract
Modern language models (LMs) are not robust to out-of-distribution inputs. Machine generated (“optimized”) prompts can be used to modulate LM outputs and induce specific behaviors while appearing completely uninterpretable. In this work, we investigate the composition of optimized prompts, as well as the mechanisms by which LMs parse and build predictions from optimized prompts. We find that optimized prompts primarily consist of punctuation and noun tokens which are more rare in the training data. Internally, optimized prompts are clearly distinguishable from natural language counterparts based on sparse subsets of the model’s activations. Across various families of instruction-tuned models, optimized prompts follow a similar path in how their representations form through the network.   [^1] Code and models available at https://github.com/rimon15/demyst_optim_prompts
@@ 1 Introduction
Language models (LMs) (Grattafiori et al., 2024; Biderman et al., 2023a; Team et al., 2024; Abdin et al., 2024) are trained on large amounts of filtered internet data (Gao et al., 2020; Raffel et al., 2020; Penedo et al., 2024; Soldaini et al., 2024), which consist primarily of interpretable natural language text. Recent work has found that these models are sensitive to machine-generated optimized prompts, which, although seemingly uninterpretable, can be used to elicit targeted behaviors (Shin et al., 2020; Wen et al., 2023; Zou et al., 2023b; Melamed et al., 2024). Specifically, we define optimized prompts as prompts that are generated via the gradient-based discrete prompt optimization method called Greedy Coordinate Gradient (GCG) (Zou et al., 2023b); see Section 2 for further background.
@@ 1.1 Our contributions
When comparing tokens in both optimized and natural language prompts to the pre-training corpus, we find that the majority of tokens in optimized prompts are more rare with respect to the training data than their natural language counterparts. Furthermore, the token distribution of optimized prompts visibly deviates from standard Zipfian behavior; see Section 5.
@@ 3 Experimental setup
We focus our work on transformer decoder (Vaswani et al., 2017; Radford and Narasimhan, 2018) models. We use the Tiny Stories Eldan and Li (2023) dataset, which consists of synthetically generated stories meant to be understandable by a three year old child. We train a transformer decoder language model based on the GPT-NeoX Black et al. (2022); Biderman et al. (2023a) architecture; see Appendix A.1 for full training details. Originally, the model uses a Byte-pair encoding (BPE) tokenizer (Sennrich et al., 2016), which results in optimized prompts that include several nonsensical characters and subwords Melamed et al. (2024); Cherepanova and Zou (2024). Because we wish to better understand which specific words appear in optimized prompts, we train a new word-level tokenizer over the Tiny Stories corpus. Using word-level tokenization allows us to better interpret optimized prompts, since we do not need to extrapolate meaning from sub-word tokens and can directly evaluate each word in the prompt individually.
@@
In addition to the word-level Tiny Stories model, we optimize prompts using 18 open models from various model families, including both base models, and their instruction-tuned variants which have been aligned for chat purposes. We use a variety of datasets for the optimization including Alpaca (Taori et al., 2023), WikiText-103 (Salesforce, 2021), OpenHermes-2.5 (Teknium, 2023), and Dolly-15k (Conover et al., 2023); see Appendix A.2 for further details.
@@
$d_{KL}({\boldsymbol{p}}^{*}||{\boldsymbol{p}})=\frac{1}{n}\sum_{i=1}^{n}\log(\mathbb{P}_{\mathrm{LM}}({\boldsymbol{d}}_{i}|{\boldsymbol{p}}^{*}))\\ -\log(\mathbb{P}_{\mathrm{LM}}({\boldsymbol{d}}_{i}|{\boldsymbol{p}})).$ (1)
@@ 4 Optimized prompts consist of specific influential tokens
$s_{i}=d_{KL}({\boldsymbol{p}}||{\boldsymbol{p}}_{-i}),$ (2)
where ${\boldsymbol{p}}_{-i}=[p_{1},...,p_{i-1},p_{i+1},...,p_{k}]$ is the prompt with token $i$ removed, and $d_{KL}$ is defined in Equation 1. A larger influence score indicates that removing token $i$ causes a greater deviation from the functional behavior of the original prompt.
@@ 5.1 Optimized prompts do not look like natural language (distributionally)
Figure 5: Sparse linear classifier probe accuracy on top-$k$ features of the model output at each layer. We identify top features using Equation 4, and train a linear classifier to discriminate between optimized and natural language prompts. The first column compares optimized and natural language prompts. The second and third columns show a baseline comparison of natural language prompts vs. other natural language prompts, and of natural language prompts vs. random prompts, respectively. We refer the reader to Figure 10 in Appendix D for the full results on all models.
@@ 5.2 Optimized prompts rely on tokens that are rare in the training data
Figure 6: Average top-10 token prediction overlap. The overlap is computed via Equation 5. Overall, the importance of the top identified features appears to be model dependent, and it is not necessarily the case that the optimized prompts rely more heavily on these features. However, it is clear that there are certain layers which influence the output more (specifically the first and last layers); see Figure 11 in Appendix D for the full results.
@@ 6.1 Sparse probing for optimized prompts
We then train a logistic regression classifier using the top identified features with varying levels of sparsity; see Appendix A.3 for additional training details. Overall, we find that the classifier is able to discriminate ground truth and optimized examples with high accuracy, even with high levels of sparsity; see Figure 5. This is in contrast to the near-random accuracy found in the baseline evaluations which compare original natural language prompts to randomly generated prompts and other natural language prompts. It is important to note that even though for optimized prompts, models clearly contain unique representations which are easily distinguishable from natural prompts, these representations still generate functionally similar outputs.
@@ 6.2 Do optimized prompts rely on a distinct subspace?
Given that we can classify prompts based on distinct features in the activations, we test the importance of these features in eliciting desired output from optimized prompts. Prior work finds that optimized prompt are sensitive to discrete perturbations (Ishibashi et al., 2023; Melamed et al., 2024; Cherepanova and Zou, 2024), and we hypothesize that this sensitivity is also present in the model’s internal representations. In order to verify our hypothesis, we perform causal intervention and zero-out the top features identified layer-wise via Equation 4, and then measure the top-10 token overlap for each (original, optimized) prompt pair.
@@
We find that intervening on top features for both natural and optimized prompts has a pronounced effect when compared to the baseline of intervening on random features. However, surprisingly, our experiment contradicts the hypothesis. As shown in Figure 6, for the majority of layers in each model, there is not a large difference between the effect of ablations of top-$k$ features on natural language prompts versus optimized prompts. This means that although optimized prompts may be more sensitive to discrete token-level perturbations than natural language prompts, they do not necessarily share this sensitivity when evaluated from the perspective of the model’s internal representations. We do note that for some models, there are specific layers (first and final) which induce a more pronounced change on the output when top features are ablated.
@@ 6.3 How do LMs build predictions from optimized prompts?
Figure 7: Layer-wise KL divergence. The layer-wise KL divergence is computed as described in Section 6.3. We find that instruction-tuned models follow a similar path; see Figure 12 in Appendix D for the full results on all models.
@@ 7 Discussion
Our work analyzes the mechanisms and ways in which language models parse and interpret discretely optimized prompts. We find that optimized prompts consist primarily of punctuation and noun tokens which are, on average, more rare in the training data than their natural language counterparts. Through sparse probing, we are able to classify optimized prompts and their natural language counterparts with high accuracy. Furthermore, when ablating neurons from model layers, the effectiveness of optimized prompts does not drop in a significant way compared to their natural language counterparts.
One possible application of our analysis is to identify optimized “jailbreak” prompts before these prompts are even fully processed by the model. For example, one can train a simple linear classifier on a set of optimized prompts and natural language, and efficiently alert the model provider if a user is inputting suspicious optimized prompts based on the classifier at intermediate layer. Although the majority of prior work has studied optimized prompts through the lens of adversarial attacks, it is also possible that such prompts are benign in nature. It will require further investigation to differentiate benign and malicious optimized prompts.
@@
Due to computational constraints, we limit our analysis to models up to 8 billion parameters, as the discrete optimization process requires significant GPU memory and compute. Nevertheless, we test 18 different models. Future work can consider analyzing larger models and investigating trends in optimized prompts across a wider array of model sizes. This will also enable research on scaling effects for prompt optimization.
@@ A.2 Optimization and data setup
For all models, we perform the evil twins optimization procedure for 500 steps, with early stopping if $d_{KL}({\boldsymbol{p}}^{*}||{\boldsymbol{p}})\leq 5.0$. In total, we obtain 5000 unique prompts for the word-stories model, and 1200 unique prompts for the open models (from the four aforementioned prompt datasets), with 300 prompts from each dataset. We then filter all final optimized prompts such that $d_{KL}({\boldsymbol{p}}^{*}||{\boldsymbol{p}})\leq 10.0$. Table 2 displays the number of prompts that were optimized for each model after filtering. For the discrete optimization, we use one 8x A100 node.
@@
A.3 Classifier probe training
The classifier is a logistic regression trained via scikit-learn (Buitinck et al., 2013) for 100 iterations with $l_{2}$ penalty, liblinear solver, and $1e^{-4}$ convergence tolerance.
@@ Appendix D Experimental results on all models
We report the full results on the remainder of the 18 tested models for the probing classifier experiment, the feature ablation experiment, and the layer-wise KL divergence experiment. Figure 10 displays the results for all model suites on the sparse probing experiment. Figure 11 displays the results for all model suites on the intervention experiment. Figure 12 displays the results for all model suites on the layer-wise KL divergence experiment.